OpenStudy (anonymous):

-.- what does it mean when you see something like this....0.05x + 0.08y = 60; x + y = 1000 i want to know what the semi colon is for

OpenStudy (anonymous):

Its just telling you that the x + y = 1000 is another equation. Its a linear way to write a system of equations. So all the information they give you is: \[0.05x + 0.08y = 60\]\[x + y = 1000\] And they want you to solve for x and y. Are you having trouble solving it? Or can you take it from there?

OpenStudy (anonymous):

ohh ok thanks!!!